DIFFERENTIAL GEOMETRY/PDE SEMINAR

Wednesday, January 19, 2005 Padelford C-36 3:50 pm

An eta invariant of Cauchy-Riemann manifolds, and applications

Olivier Biquard (Institut de Recherche Mathmatique Avance, Strasbourg, France)

The eta invariant is a conformal invariant of 3-manifolds, relating Riemannian geometry, topology and spectral theory. I shall explain how an analogous invariant can be constructed for strictly pseudoconvex CR manifolds, following a general philosophy on relations between conformal invariants and Einstein metrics. I shall discuss applications to filling CR manifolds and relations with spectral theory.

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.